OpenStudy (anonymous):

Solve for x. 30=(x+3)+(x+3)+x.

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

constants with constants and variables with variables

OpenStudy (anonymous):

I'm a 6th gradeer so yeah

OpenStudy (anonymous):

Thanks though

OpenStudy (anonymous):

30=(x+3)+(x+3)+x The ( ) don't matter because it is addition, so we can just get rid of those 30=x+3+x+3+x

OpenStudy (anonymous):

So its 30=(x+3)+(x+3)+x -----> 30=4x+4x+x ------->30= 9x -------> x=3.3

OpenStudy (anonymous):

30=x+3+x+3+x Combining like terms 30 = 3x + 6

OpenStudy (anonymous):

You are allowed to subtract, add, multiply, divide anything from both sides to simplify the equation. We want to get x by itself. 30 - 6 = 3x + 6 - 6 24 = 3x divide both sides by 3 x = 8

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (anonymous):

Know i can do the rest :D

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

NOw*

OpenStudy (anonymous):

:D